The Framework of a wall is 3 1/2 in. thick. We apply 5/8 in. wallboard and 1/4 in. paneling to the inside. Siding that is 3/4 in. thick is applied to the outside. What is the finished thickness of the wall?
Answer by scott8148(6628)

3 1/2 + 5/8 + 1/4 + 3/4 = 3 4/8 + 5/8 + 1 = 4 1/8 + 1 = 5 1/8